Grade 11Mechanics

The orbital velocity of an artificial satellite in a circular orbit just above the centre’s surface is v.
For a satellite orbiting at an altitude of half of the earth’s radius, the orbital velocity is
(A) 3v/2 (B) 3 / 2v
(C) 2 / 3 v (D) (2/3)v

DEBASISH
8 Years agoGrade 11

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear student
 
As we know
Vo = \sqrt(Gm/(R +h))
Vo \alpha  1/ \sqrt(R+h)
Hence
V2/V1 = \sqrt[(R+0) /(R +R/2)]
V2 = \sqrt(2/3) * V
 
Regards
Arun (askIITians forum expert)